OpenStudy (anonymous):

Which of the following represents the area of a rectangle whose length is 3x – 5 and whose width is x – 4? 3x2 + 17x + 20 3x2 – 17x + 20 3x2 – 20 3x2 + 20

OpenStudy (anonymous):

The area of a rectangle is length times width. (3x - 5)(x - 4) = 3x^2 -5x -12x + 20 = 3x^2 -17x +20 = Area